OpenStudy (devonhoward15):

Which counterexample disproves the conjecture? If n is a positive integer, then 2^n-1 is a prime number

OpenStudy (devonhoward15):

n=6 5 3 2

OpenStudy (chrishogan):

are the rest of those like A. n=6 B. n=5 C. n=3 D. n=2?

OpenStudy (devonhoward15):

yes... i did it the short way srry lol

OpenStudy (chrishogan):

Aight B is a counter example because 2(5) - 1 = 10 - 1 = 9 is NOT a prime number.

OpenStudy (welshfella):

replace n by the given choices and see what the values are

OpenStudy (chrishogan):

So n=5 would be your answer.

OpenStudy (welshfella):

No Chris its 2 to the power n not 2 times n.

OpenStudy (welshfella):

last one 2 is 2^2 - 1 = 4-1 = 3 which is a prime number so 2 is not a counterexample.

OpenStudy (chrishogan):

Opp my mistake, you right.